# Independence and Induction in Reverse Mathematics

### v2

@@ Abstract.
We continue the project of the study of reverse mathematics principles inspired by cardinal invariants. In this article in particular we focus on principles encapsulating the existence of large families of objects that are in some sense mutually independent. More precisely, we study the principle $\mathsf{MAD}$ stating that a maximal family of pairwise almost disjoint sets exists; and the principle $\mathsf{MED}$ expressing the existence of a maximal family of functions that are pairwise eventually different. We investigate characterisations of and relations between these principles and some of their variants. It will turn out that induction strength is an essential parameter in this context.
@@ 1. Introduction
Reverse mathematics is the program of classifying the strength of mathematical statements relative to each other over a base system—see Simpson [6]. These mathematical statements are often referred to as “principles,” as in the Ramsey-style “combinatorial principles” that have captured such attention since the mid-1990’s.
@@ 2. Preliminaries
We are interested in models in the language of second-order arithmetic. These take the form $\mathcal{M}=(M,\mathcal{S},+,\cdot,0,1,\in)$ where $M$ is the first-order part and $\mathcal{S}$ is a subset of the power set of $M$. Typically we suppress the operations and write simply $\mathcal{M}=(M,\mathcal{S})$. We call the elements of $M$ numbers, with the idea that they stand in for the natural numbers $\omega$ and obey many of the same axioms. In this article we resist the convention of writing “sets” to mean the elements of $\mathcal{S}$, because we will often need to discuss sets $A\subseteq M$ which are possibly outside of $\mathcal{S}$. Call a formula a $\Sigma^{A}_{n}$ formula if it is $\Sigma_{n}$ while using the given set $A$ as a second-order parameter (which we also call an oracle). Call a set $B\subseteq M$ a $\Sigma^{A}_{n}$ set if $B$ is definable by a $\Sigma_{n}^{A}$ formula. Formulas with multiple oracles are allowed; call a formula or set $\Sigma^{0}_{n}$ if it is $\Sigma^{A_{1},\ldots,A_{k}}_{n}$ for some $A_{1},\ldots,A_{k}\in\mathcal{S}$; this includes the case where $k=0$ and the list is empty. Make all the analogous definitions with $\Pi$ and $\Delta$ in place of $\Sigma$.
Here are some of the axioms we will consider. Let $P^{-}$ be the axioms of Peano arithmetic not including the induction scheme. Let $\mathsf{I}\mathrm{\Sigma}_{n}^{0}$ denote the induction scheme for $\Sigma^{0}_{n}$ formulas, and let $\mathsf{B}\mathrm{\Sigma}_{n}^{0}$ denote the $\Sigma^{0}_{n}$-bounding scheme. Kirby and Paris [3] proved over the system $P^{-}$ that $\mathsf{B}\mathrm{\Sigma}_{n+1}^{0}\rightarrow\mathsf{I}\mathrm{\Sigma}_{n}^{0}$ and $\mathsf{I}\mathrm{\Sigma}_{n}^{0}\rightarrow\mathsf{B}\mathrm{\Sigma}_{n}^{0}$ for $n\geq 1$, and that these implications are strict. (In their account there were no second-order parameters, but the proof is the same.) The system $\mathsf{RCA}_{0}$ consists of $P^{-}$ together with $\mathsf{I}\mathrm{\Sigma}_{1}^{0}$ and the comprehension scheme for $\Delta^{0}_{1}$ sets. In particular, if $\mathcal{M}=(M,\mathcal{S})$ is a model of $\mathsf{RCA}_{0}$, then $S$ is closed under Turing reducibility.
@@
By identifying a set $A$ with its characteristic function, we can naturally speak of a family ${\mathcal{F}}$ of sets $A\subseteq M$ being weakly represented. If ${\mathcal{F}}$ is a weakly represented family (of functions or of sets), then each element $f\in{\mathcal{F}}$ is $\Delta^{0}_{1}$ relative to sets in $S$ and therefore is in $\mathcal{S}$. By considering the model $\mathcal{M}$’s version of the universal oracle Turing machine, we obtain the following result.
@@
The rest of this paper is structured as follows. Section 3 deals with the principle $\mathsf{DOM}$ which expresses essentially that there is a function dominating every total function in a weakly represented family; we prove a version of Martin’s theorem [5] characterizing such functions in terms of high Turing degrees. Section 4 is a technical section where we carry out a gener al $0^{\prime\prime}$ tree argument for use in later sections, to construct new weakly represented families. Section 5 deals with the principle $\mathsf{MAD}$, which says that there is a maximal almost disjoint (MAD) family of sets. We obtain results comparing $\mathsf{MAD}$ with $\mathsf{DOM}$, and with $\mathsf{I}\mathrm{\Sigma}_{2}^{0}$. In Section 6, we investigate the principles $\mathsf{MED}$, which says there is a maximal eventually different (MED) family of functions, as well as $\mathsf{AVOID}$, which says essentially that there is a function eventually different (ED) from all total functions in a weakly represented family; we obtain results about the relationship between the two. Section 7 covers the principle $\mathsf{NDF}$, which says essentially that there is a hyperimmune Turing degree. Finally, Section 8 is concerned with the principle $\mathsf{BI}$, which says that there is a bi-immune Turing degree; we use the existence of a low bi-immune Turing degree to obtain a conservation result.
@@ 3. Dominating families of functions
In an $\omega$-model—meaning a model $(M,\mathcal{S})$ in which $M$ is the true natural numbers—if $X\in\mathcal{S}$ computes a function which dominates every total recursive function, then $X$ is generalized high, that is, $X^{\prime}\equiv_{T}(X\oplus\emptyset^{\prime})^{\prime}$. This was first proved by Martin [5]. The phenomenon is reflected in $\mathsf{RCA}_{0}$ by the following proposition, proved essentially in Chong, Qian, Slaman and Yang [1, Lemma 2]:
Proposition 5 (Martin’s Theorem, formalized).
Over $\mathsf{RCA}_{0}$, the following are equivalent:
@@
    $\mathsf{DOM}$
@@
    $\mathsf{HIGH}{:}$ For every $A$, there is a $B$ such that every $\Sigma^{A}_{2}$ is $\Delta^{B}_{2}$.
@@
    For each $A$, every $\Sigma^{A}_{2}$ set is $\Delta^{B}_{2}$ for some $B$.
@@
This formalized version of Martin’s Theorem allows for a short proof of the previously known fact that $\mathsf{DOM}$ and $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ together imply full arithmetical induction.
@@
$\mathsf{RCA}_{0}+\mathsf{DOM}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ implies $\mathsf{I}\mathrm{\Sigma}_{n}^{0}$ for every $n\geq 1$.
@@
The following holds for every model $(M,\mathcal{S})$ of $\mathsf{RCA}_{0}$:
@@
Fix $(M,\mathcal{S})\models\text{RCA}_{0}+\mathsf{DOM}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ and let $A_{0}\in\mathcal{S}$. Use Proposition 5 to obtain a sequence $\{A_{n}:n\in\omega\}$ such that $A_{n+1}$ is high relative to $A_{n}$. In particular, the double jump $A_{n}^{\prime\prime}$ is $\Delta_{2}^{A_{n+1}}$ and therefore regular by $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ and Lemma 7. Since $A_{0}^{\prime\prime}\leq_{T}A_{1}^{\prime}$, we also have $A_{0}^{\prime\prime\prime}\leq_{T}A_{1}^{\prime\prime}$. Hence by transitivity ($\mathsf{B}\mathrm{\Sigma}_{1}$ relative to $A_{2}^{\prime}$), we have $A_{0}^{\prime\prime\prime}\leq_{T}A_{2}^{\prime}$ and hence that $A_{0}^{\prime\prime\prime}$ is regular. Iterating this argument for larger and larger $n$, we find that $A_{0}^{(n)}\leq_{T}A_{n-1}^{\prime}$ and $A_{0}^{(n)}$ is regular for all $n\in\omega$. The result follows from Lemma 7 and the fact that $A_{0}$ was arbitrary. ∎
@@ 4. Handling index sets
$\{\Psi_{e_{i}}\colon e_{i}\leavevmode\nobreak\ is\leavevmode\nobreak\ least\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ \varphi(\langle e_{0},\ldots,e_{i-1}\rangle)\textnormal{ holds}\}$
@@
5. Maximal, almost disjoint families of sets
@@
Hölzl, Raghavan, Stephan and Zhang showed in [2] that $\mathsf{MAD}$ is equivalent to $\neg\mathsf{DOM}$ in $\omega$-models. We extend this result to all models of $\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}$.
@@
=${\mathcal{F}}=\{\Phi^{A}_{e_{i}}\colon e_{i}\in\langle e_{0},e_{1},\ldots\}$
, where $\langle e_{0},e_{1},\ldots\rangle$ is lexicographically least satisfying both $\psi$ and the relation $e_{0}<e_{1}<\cdots$.
@@
An alternate proof of $\neg\mathsf{MAD}\rightarrow\mathsf{DOM}$ might construct the partition directly using a $0^{\prime\prime}$ tree argument similar to the proof of Lemma 10. A natural next question is whether $\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ is necessary for the equivalence $\neg\mathsf{MAD}\leftrightarrow\mathsf{DOM}$ to hold. The answer is yes, although this is in part an artefact of our choice of definition for finite weakly represented families.
@@
Then $\Psi^{A}$ weakly represents the family =${\mathcal{F}}=\{C_{i}\colon i\in B\}$. Since the $C_{i}$ form a partition, ${\mathcal{F}}$ is almost disjoint; we claim that ${\mathcal{F}}$ is MAD. To see this, suppose for a contradiction that some $D\in\mathcal{S}$ is $\mathcal{M}$-infinite, but $D\cap C_{i}$ is $\mathcal{M}$-finite for every $i\in B$. Then $D\cap C_{i}$ is $\mathcal{M}$-finite for every $i<a$. That is, we have $(\forall i<a)(\exists s)(\forall t>s)[t\not\in C_{i}\cap D]$. Using $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ to exchange the two leading quantifiers yields $(\exists s)(\forall i<a)(\forall t>s)[t\not\in C_{i}\cap D]$. Since the $C_{i}$ form a partition, this implies that $D$ is $\mathcal{M}$-finite, a contradiction.
@@ 6. Avoidance and eventual difference
Statement 16 (Avoidance principle, $\mathsf{AVOID}$).
@@
Statement 17 (Maximal Eventually Different Family principle, $\mathsf{MED}$).
@@
The following is a reformulation of a previously known result in the terminology of reverse mathematics; we include our own proof since as far as we know it is more direct than others that have appeared in print.
Proposition 19 (Kjos-Hanssen, Merkle, Stefan [4]).
@@
Suppose $\mathcal{M}\models\mathsf{RCA}_{0}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}+\neg\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ is topped by $A$. Let $\Phi^{A}$ be universal in the sense of Lemma 2, suppose $c\in M$ is large enough that $\{e<c:\Phi_{e}^{A}$ is not $\mathcal{M}$-finite}}\}}, and define a partial function $u(e,x)$ by:
@@
$(\forall i<c)(\exists d)(\forall x>d)[\textnormal{either\leavevmode\nobreak\ }u(i,x){\uparrow}\textnormal{\leavevmode\nobreak\ or\leavevmode\nobreak\ }u(i,x)<u(e,x)]$
Using $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ to exchange the first two quantifiers, this becomes:
$(\exists d)(\forall i<c)(\forall x>d)[\textnormal{either\leavevmode\nobreak\ }u(i,x){\uparrow}\textnormal{\leavevmode\nobreak\ or\leavevmode\nobreak\ }u(i,x)<u(e,x)]$
But this means, with knowledge of $d$ and $e$, we the set $\{i<c\colon\Phi_{e}^{A}$ is total}}\}} is $\Pi^{0}_{1}$ and hence $\mathcal{M}$-finite, a contradiction. ∎
@@
We define by stages a set $(g_{i})_{i<c}$ of partial functions; at stage $s=u(e_{i},x_{i})$, we define $g_{i}(x_{i})$. Begin by forming the set $E=\{e<c\colon\Phi^{A}_{e,s}{\downarrow}$ and $(\forall j\neq i)g_{j}(x)\neq\Phi^{A}_{e,s}(x)\}$. If $E$ is empty, set $g_{i}(x)$ to equal some number that has not been seen previously in the construction. Otherwise, select an $e$ which minimizes $e+|\{y<x\colon\Phi_{e,s}(y)=g_{i}(y)\}|$. This completes the constrution.
We make four claims about $(g_{i})_{i<c}$. The first is that if $i\neq j$, then $g_{i}(x)\neq g_{j}(x)$ for all $x$ which are both defined; this is direct from the construction. The second is that those $g_{i}$ which are total are eventually different; this claim follows a fortiori from the first. The third is that $g_{i}$ is total iff the domain of $g_{i}$ is unbounded in $M$ iff $\Phi_{e}^{A}$ is total; this is direct from the construction and the definition of $u(i,x)$. The fourth claim is that these $g_{i}$ form a maximal such family. For a contradiction, suppose the fourth claim is false, which since the model is topped means that there is some $\Phi^{A}_{e}$ which is total and eventually different from each total $g_{i}$. In symbols:
@@
Using $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ to exchange the leading quantifiers, this becomes:
@@
Keeping $e$ fixed, let $i$ be as provided by the Lemma. Consider the number $n(x)=e+|\{y<x\colon\Phi^{A}_{e,s}(y)=g_{i}(y)\}|$ which was used in the construction of $g_{i}$. Clearly $n$ is nondecreasing; the equation above tells us is that $n(x)$ reaches some $\mathcal{M}$-finite limit $N$, after some stage $s=\langle i,x\rangle$ of the construction. Then at each stage $t=\langle i,x^{\prime}\rangle$ which is greater than $s$, we know that $\Phi_{e}^{A}(x^{\prime})$ was not chosen to equal $g_{i}(x^{\prime})$. Since $i$ is given by the Lemma, we know that $u(e,x)<u(i,x)$ for $\mathcal{M}$-infinitely many $x$; hence at $\mathcal{M}$-infinitely many stages, there was another $e^{\prime}\in E$ with
$e^{\prime}+|\{y<x\colon\Phi_{e^{\prime},s}(y)=g_{i}(y)\}|<N$
But this is possible only if $e^{\prime}<N$, and then no more than $N-e^{\prime}$ times. This contradicts the pigeonhole principle, and proves the fourth claim. Therefore $(g_{i})_{i<c}$ is a MED family, as required. ∎
7. Nondominated families of functions
The following principle is analogous to the existences of hyperimmune degree.
Statement 22 (Nondominated Family principle, $\mathsf{NDF}$).
Given any weakly represented family $\mathcal{F}$ of functions, there exists a function $g$ such that for all $f\in\mathcal{F}$, there are unboundedly many $x\in M$ such that $g(x)>f(x)$; that is, no $f\in\mathcal{F}$ dominates $g$.

### v3

@@ Abstract.
We continue the project of the study of reverse mathematics principles inspired by cardinal invariants. In this article in particular we focus on principles encapsulating the existence of large families of objects that are in some sense mutually independent. More precisely, we study the principle $\mathsf{MAD}$ stating that a maximal family of pairwise almost disjoint sets exists; and the principle $\mathsf{MED}$ expressing the existence of a maximal family of functions that are pairwise eventually different. We investigate characterisations of and relations between these principles and some of their variants. It turns out that induction strength at the levels of $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ or $\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ is an essential parameter; for instance, over $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$, we show that $\neg\mathsf{MAD}$ is equivalent to the principle $\mathsf{DOM}$ expressing that every weakly represented family of functions is dominated by some other function.
@@ 1. Introduction
Reverse mathematics is the program of classifying the strength of mathematical statements relative to each other over a base system — see Simpson [8]. These mathematical statements are often referred to as “principles,” as in the Ramsey-style “combinatorial principles” that have captured much attention since the mid-1990’s.
@@ 2. Preliminaries
We are interested in models in the language of second-order arithmetic. These take the form $\mathcal{M}=(M,\mathcal{S},+,\cdot,0,1,\in)$ where $M$ is the first-order part and $\mathcal{S}$ is a subset of the power set of $M$. Typically we suppress the operations and write simply $\mathcal{M}=(M,\mathcal{S})$. We call the elements of $M$ numbers, with the idea that they stand in for the natural numbers $\omega$ and obey many of the same axioms. In this article we resist the convention of writing “sets” to mean the elements of $\mathcal{S}$, because we will often need to discuss sets $A\subseteq M$ which are possibly outside of $\mathcal{S}$. Call a formula a $\Sigma^{A}_{n}$ formula if it is $\Sigma_{n}$ while using the given set $A$ as a second-order parameter (which we also call an oracle). Call a set $B\subseteq M$ a $\Sigma^{A}_{n}$ set if $B$ is definable by a $\Sigma_{n}^{A}$ formula. Formulas with multiple oracles are allowed; call a formula or set $\Sigma^{0}_{n}$ if it is $\Sigma^{A_{1},\ldots,A_{k}}_{n}$ for some $A_{1},\ldots,A_{k}\in\mathcal{S}$; this includes the case where $k=0$ and the list is empty. Analogous definitions with “$\Pi$” or “$\Delta$” in place of “$\Sigma$” can be given in the obvious ways.
Here are some of the axioms we will consider. Let $P^{-}$ be the axioms of Peano arithmetic not including any induction schemes. Let $\mathsf{I}\mathrm{\Sigma}_{n}^{0}$ denote the induction scheme for $\Sigma^{0}_{n}$ formulas, and let $\mathsf{B}\mathrm{\Sigma}_{n}^{0}$ denote the $\Sigma^{0}_{n}$-bounding scheme. Over $P^{-}$, Kirby and Paris [4] proved the following strict implications for every $n\geq 1$:
$\mathsf{B}\mathrm{\Sigma}_{n+1}^{0}\rightarrow\mathsf{I}\mathrm{\Sigma}_{n}^{0}\quad\text{and}\quad\mathsf{I}\mathrm{\Sigma}_{n}^{0}\rightarrow\mathsf{B}\mathrm{\Sigma}_{n}^{0}$
(In their account there were no second-order parameters, but the proof is the same.) The system $\mathsf{RCA}_{0}$ consists of $P^{-}$ together with $\mathsf{I}\mathrm{\Sigma}_{1}^{0}$ and the comprehension scheme for $\Delta^{0}_{1}$ sets. In particular, if $\mathcal{M}=(M,\mathcal{S})$ is a model of $\mathsf{RCA}_{0}$, then $S$ is closed under Turing reducibility.
@@
By identifying a set $A$ with its characteristic function, we can naturally speak of a family ${\mathcal{F}}$ of sets $A\subseteq M$ being weakly represented. We mention that these notions of weakly represented families also were used by Hölzl, Jain and Stephan [2] to study inductive inference in the setting of reverse mathematics.
If ${\mathcal{F}}$ is a weakly represented family (of functions or of sets), then each element $f\in{\mathcal{F}}$ is $\Delta^{0}_{1}$ relative to sets in $S$ and therefore is in $\mathcal{S}$. By considering the model $\mathcal{M}$’s version of the universal oracle Turing machine, we obtain the following result.
@@
The rest of this paper is structured as follows. Section 3 deals with the principle $\mathsf{DOM}$ which expresses essentially that there is a function dominating every total function in a weakly represented family; we prove a version of Martin’s theorem [6] characterizing such functions in terms of high Turing degrees. Section 4 is a technical section where we carry out a gener al $0^{\prime\prime}$ tree argument for use in later sections, to construct new weakly represented families. Section 5 deals with the principle $\mathsf{MAD}$, which says that there is a maximal almost disjoint (MAD) family of sets. We obtain results comparing $\mathsf{MAD}$ with $\mathsf{DOM}$, and with $\mathsf{I}\mathrm{\Sigma}_{2}^{0}$. In Section 6, we investigate the principles $\mathsf{MED}$, which says there is a maximal eventually different (MED) family of functions, as well as $\mathsf{AVOID}$, which says essentially that for every weakly represented family there is a function eventually different (ED) from all functions in the family; we obtain results about the relationship between the two. Section 7 covers the principle $\mathsf{HI}$, which says essentially that there is a hyperimmune Turing degree. Finally, Section 8 is concerned with the principle $\mathsf{BI}$, which says that there is a bi-immune Turing degree; we use the existence of a low bi-immune Turing degree to obtain a conservation result.
@@ 3. Dominating families of functions
In an $\omega$-model — meaning a model $(M,\mathcal{S})$ in which $M$ is the true natural numbers — if $X\in\mathcal{S}$ computes a function which dominates every total recursive function, then $X$ is generalised high, that is, $X^{\prime}\equiv_{T}(X\oplus\emptyset^{\prime})^{\prime}$. This was first proven by Martin [6]. The phenomenon is reflected in $\mathsf{RCA}_{0}$ by the following proposition, proven essentially by Chong, Qian, Slaman and Yang [1, Lemma 2]:
Proposition 5 (Martin’s Theorem, formalised).
For a model $(M,\mathcal{S})$, the following are equivalent over $\mathsf{RCA}_{0}$:
@@
    $\mathsf{DOM}{:}$ For every $\mathcal{F}$ that is a weakly represented family in $\mathcal{S}$ there is a function $g\in\mathcal{S}$ which grows faster than every $f\in\mathcal{F}$.
@@
    $\mathsf{HIGH}{:}$ For every $A\in\mathcal{S}$ there is a $B\in\mathcal{S}$ such that every $\Sigma^{A}_{2}$ set $C\subseteq M$ is $\Delta^{B}_{2}$.
@@
    For each $A\in\mathcal{S}$ and every $\Sigma^{A}_{2}$ set $C\subseteq M$ there is a $B\in\mathcal{S}$ such that $C$ is $\Delta^{B}_{2}$.
Note that we do not assume $C\in\mathcal{S}$.
@@
This formalised version of Martin’s Theorem allows for a short proof of the previously known fact that $\mathsf{DOM}$ and $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ together imply full arithmetical induction and thus Peano Arithmetic $(\mathsf{PA})$.
@@
$\mathsf{RCA}_{0}+\mathsf{DOM}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ implies $\mathsf{I}\mathrm{\Sigma}_{n}^{0}$ for every $n\geq 1$. In short, $\mathsf{RCA}_{0}+\mathsf{DOM}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}\vdash\mathsf{PA}$.
@@
The following statements hold for every model $(M,\mathcal{S})$ of $\mathsf{RCA}_{0}$:
@@
Fix $(M,\mathcal{S})\models\text{RCA}_{0}+\mathsf{DOM}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ and let $A_{0}\in\mathcal{S}$. Use Proposition 5 to obtain a sequence $\{A_{n}\colon n\in\omega\}$ such that $A_{n+1}$ is high relative to $A_{n}$. In particular, the double jump $A_{n}^{\prime\prime}$ is $\Delta_{2}^{A_{n+1}}$ and therefore regular by $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ and Lemma 7. Since $A_{0}^{\prime\prime}\leq_{T}A_{1}^{\prime}$, we also have $A_{0}^{\prime\prime\prime}\leq_{T}A_{1}^{\prime\prime}$. Hence by transitivity ($\mathsf{B}\mathrm{\Sigma}_{1}$ relative to $A_{2}^{\prime}$), we have $A_{0}^{\prime\prime\prime}\leq_{T}A_{2}^{\prime}$ and hence that $A_{0}^{\prime\prime\prime}$ is regular. Iterating this argument for larger and larger $n$, we find that $A_{0}^{(n)}\leq_{T}A_{n-1}^{\prime}$ and $A_{0}^{(n)}$ is regular for all $n\in\omega$. The result follows from Lemma 7 and the fact that $A_{0}$ was arbitrary. ∎
@@ 4. Handling index sets
$\{\Psi_{e_{i}}\colon e_{i}\textnormal{ is least such that }\varphi(\langle e_{0},\ldots,e_{i-1}\rangle)\textnormal{ holds}\}$
@@
5. Maximal almost disjoint families of sets
@@
Hölzl, Raghavan, Stephan and Zhang [3] showed that $\mathsf{MAD}$ is equivalent to $\neg\mathsf{DOM}$ in $\omega$-models. We extend this result to all models of $\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}$.
@@
=${\mathcal{F}}=\{\Phi^{A}_{e_{i}}\colon e_{i}\in\langle e_{0},e_{1},\ldots\rangle\},$
where $\langle e_{0},e_{1},\ldots\rangle$ is lexicographically least satisfying both $\psi$ and the relation $e_{0}<e_{1}<\cdots$.
@@
An alternate proof of $\neg\mathsf{MAD}\rightarrow\mathsf{DOM}$ might construct the partition directly using a $0^{\prime\prime}$ tree argument similar to the proof of Lemma 10.
@@
Then $\Psi^{A}$ weakly represents the family =${\mathcal{F}}=\{C_{i}\colon i\in B\}$. Since the $C_{i}$ form a partition, ${\mathcal{F}}$ is almost disjoint; we claim that ${\mathcal{F}}$ is MAD. To see this, suppose for a contradiction that some $D\in\mathcal{S}$ is $\mathcal{M}$-infinite, but ${D\cap C_{i}}$ is $\mathcal{M}$-finite for every $i\in B$. Then $D\cap C_{i}$ is $\mathcal{M}$-finite for every $i<a$. That is, we have
$(\forall i<a)(\exists s)(\forall t>s)[t\not\in C_{i}\cap D].$
Using $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ to exchange the two leading quantifiers yields $(\exists s)(\forall i<a)(\forall t>s)[t\not\in C_{i}\cap D]$. Since the $C_{i}$ form a partition, this implies that $D$ is $\mathcal{M}$-finite, a contradiction.
@@
As a corollary, we obtain an extension of Theorem 14.
Corollary 16.
$\mathsf{RCA}_{0}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}\vdash\mathsf{DOM}\leftrightarrow\neg\mathsf{MAD}$.
Proof.
As $\mathsf{RCA}_{0}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}+\mathsf{DOM}\vdash\mathsf{I}\mathrm{\Sigma}_{2}^{0}$, no model of $\mathsf{RCA}_{0}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}+\neg\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ satisfies $\mathsf{DOM}$. But by Theorem 14 each such model satisfies $\mathsf{MAD}$. ∎
@@ 6. Avoidance and eventual difference
Statement 17 (Avoidance principle, $\mathsf{AVOID}$).
@@
Statement 18 (Maximal Eventually Different Family principle, $\mathsf{MED}$).
@@
The following proposition is a reformulation of a previously known result in the terminology of reverse mathematics; we include our own proof since as far as we know it is more direct than others that have appeared in print.
Proposition 20 (Kjos-Hanssen, Merkle, Stephan [5]).
@@
Suppose $\mathcal{M}\models\mathsf{RCA}_{0}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}+\neg\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ is topped by $A$. Let $\Phi^{A}$ be universal in the sense of Lemma 2, suppose $c\in M$ is large enough that $\{e<c\colon\Phi_{e}^{A}$ is not $\mathcal{M}$-finite}}\}}, and define a partial function $u(e,x)$ via
@@
$(\forall i<c)(\exists d)(\forall x>d)[\textnormal{either\leavevmode\nobreak\ }u(i,x){\uparrow}\textnormal{\leavevmode\nobreak\ or\leavevmode\nobreak\ }u(i,x)<u(e,x)].$
Using $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ to exchange the first two quantifiers, it follows that
$(\exists d)(\forall i<c)(\forall x>d)[\textnormal{either\leavevmode\nobreak\ }u(i,x){\uparrow}\textnormal{\leavevmode\nobreak\ or\leavevmode\nobreak\ }u(i,x)<u(e,x)].$
But this means, with knowledge of $d$ and $e$, we see that the set $\{i<c\colon\Phi_{i}^{A}$ is total}}\}} is $\Pi^{0}_{1}$ and hence $\mathcal{M}$-finite, a contradiction. ∎
The next proposition shows that for topped models satisfying $\mathsf{B}\mathrm{\Sigma}_{2}^{0}+\neg\mathsf{I}\mathrm{\Sigma}_{2}^{0}$, the axiom $\neg\mathsf{AVOID}$ coincides with $\mathsf{MED}$. Topped models satisfy that there is an oracle $A$ in the second order part such that every function is $A$-recursive. Then the $A$-recursive functions form a weakly represented family and thus all functions are in this family, thus this family witnesses $\neg\mathsf{AVOID}$. This motivates the following proposition which shows that under $\mathsf{RCA}_{0}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}+\neg\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ topped models satisfy $\mathsf{MED}$.
@@
We define by stages a set $(g_{i})_{i<c}$ of partial functions; at stage $s=u(e_{i},x_{i})$, we define $g_{i}(x_{i})$. Begin by forming the set $E=\{e<c\colon\Phi^{A}_{e,s}{\downarrow}$ and $(\forall j\neq i)g_{j}(x)\neq\Phi^{A}_{e,s}(x)\}$. If $E$ is empty, set $g_{i}(x)$ to equal some number that has not been seen previously in the construction. Otherwise, select an $e$ which minimises ${e+|\{y<x\colon\Phi_{e,s}(y)=g_{i}(y)\}|}$. This completes the constrution.
We make four claims about $(g_{i})_{i<c}$. The first is that if $i\neq j$, then $g_{i}(x)\neq g_{j}(x)$ for all $x$ which are both defined; this is direct from the construction. The second is that those $g_{i}$ which are total are eventually different; this claim follows a fortiori from the first. The third is that $g_{i}$ is total iff the domain of $g_{i}$ is unbounded in $M$ iff $\Phi_{e}^{A}$ is total; this is direct from the construction and the definition of $u(i,x)$. The fourth claim is that these $g_{i}$ form a maximal such family. For a contradiction, suppose the fourth claim is false, which since the model is topped means that there is some $\Phi^{A}_{e}$ which is total and eventually different from each total $g_{i}$. That is, formally,
@@
Using $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ to exchange the leading quantifiers, this becomes
@@
Keeping $e$ fixed, let $i$ be as provided by the Lemma. Consider ${n(x)=e+|\{y<x\colon\Phi^{A}_{e,s}(y)=g_{i}(y)\}|}$, the function used during the construction of $g_{i}$. Clearly $n$ is nondecreasing; and the equation above tells us is that $n(x)$ reaches some $\mathcal{M}$-finite limit $N$, after some stage $s=\langle i,x\rangle$ of the construction. Then at each stage ${t=\langle i,x^{\prime}\rangle}$ that is greater than $s$, we know that $\Phi_{e}^{A}(x^{\prime})$ was not chosen to equal $g_{i}(x^{\prime})$. Since $i$ is given by the Lemma, we know that $u(e,x)<u(i,x)$ for $\mathcal{M}$-infinitely many $x$; hence at $\mathcal{M}$-infinitely many stages, there was another $e^{\prime}\in E$ with
$e^{\prime}+|\{y<x\colon\Phi_{e^{\prime},s}(y)=g_{i}(y)\}|<N.$
But this is possible only if $e^{\prime}<N$, and thus no more than $N-e^{\prime}$ times. This contradicts the pigeonhole principle, and proves the fourth claim. Therefore $(g_{i})_{i<c}$ is a MED family, as required. ∎
7. Hyperimmunity in Reverse Mathematics
The following principle is a reverse mathematics formalisation of the existences of hyperimmune degrees using weakly represented families.
Statement 23 (Hyperimmunity principle, $\mathsf{HI}$).
Given any weakly represented family $\mathcal{F}$ of functions, there exists a function $g$ such that for all $f\in\mathcal{F}$, there are unboundedly many $x\in M$ such that $g(x)>f(x)$; in other words, there is no $f\in\mathcal{F}$ such that $g$ is almost everywhere less than $f$.